At their closest approach, venus and earth are 4.20*10^10 m apart. The mass of Venus is 4.87*10^24 , the mass of the earth is 5.97*10^24 kg
and G=6.67*10^-11 N.m^2/kg^2
what is the force of gravity between the two planets
please write clearly :)
Answer #1

We know that the force of the gravity

F=GMeMv/R2

Here Me =5.97*1024 kg , Mv =4.87*1024 , R=4.20*1010 m ,G=6.67*10^-11 N.m^2/kg^2

Put these values in above equation

F=6.67*10^-11*5.97*1024*4.87*1024/(4.2*1010)2

F=10.99*1017 N

F=1.10*1018 N Answer